Suppose Mike invests $800 in an account that gains 0.09% interest and is compounded quarterly. How much interest would he make in 3 years?
Answer by math_tutor2020(3817)


Use the compound interest formula
A = P*(1+r/n)^(n*t)
where,
A = final amount
P = initial amount or deposit = 800 dollars
r = annual interest rate = 0.0009
n = compounding frequency = 4
t = number of years = 3
Note: 0.09% = 0.09/100 = 0.0009

So,
A = P*(1+r/n)^(n*t)
A = 800*(1+0.0009/4)^(4*3)
A = 802.162675005764
A = 802.16
is the value of the account after 3 years.

The amount of interest earned is
A - P = 802.16 - 800 = 2.16

Answer: $2.16